solve....5.3x-2.8=4.24-3.5x.....x=?
Answer by EricL(29)

Hello!
To solve this problem, 5.3x-2.8=4.24-3.5x
Move all terms not containing x to the right-hand side of the equation.
8.8x-2.8=4.24
Move all terms not containing x to the right-hand side of the equation.
8.8x=7.04
Divide each term in the equation by 8.8.
x=0.8
--------------------
If you have any questions, feel free to ask.
--------------------
- Eric L :)